# This is the title

### v2

@@ 1. Introduction
Let $\mathcal{H}$ be a complex Hilbert space and $A$ be a possibly unbounded self-adjoint linear operator defined on domain $\mathcal{D}(A)\subseteq\mathcal{H}$. For $h\in\mathcal{D}(A)$ with $\|h\|=1$, define the uncertainty of $A$ at the point $h$ as
@@
In 1929, Robertson [7] derived the following mathematical form of the uncertainty principle (also known as uncertainty relation) of Heisenberg derived in 1927 [4] (English translation of 1927 original article by Heisenberg). Recall that for two linear operators $A:\mathcal{D}(A)\to\mathcal{H}$ and $B:\mathcal{D}(B)\to\mathcal{H}$, we define $[A,B]\coloneqq AB-BA$ and $\{A,B\}\coloneqq AB+BA$.
Theorem 1.1.
@@
In this short note, we answer Question 1.3 by deriving an uncertainty principle for Lipschitz maps acting on subsets of Banach spaces. We note that there is a Banach space version of uncertainty principle by Goh and Goodman [3] which differs from the results in this paper. We also note that nonlinear Maccone-Pati uncertainty principle is derived in [5]. It is interesting to note that the uncertainty principle of Game theory, derived by Székely and Rizzo is nonlinear [9].
@@ 2. Nonlinear Heisenberg-Robertson-Schrodinger Uncertainty Principle
Let $\mathcal{X}$ be a Banach space. Recall that the collection of all Lipschitz functions $f:\mathcal{X}\to\mathbb{C}$ satisfying $f(0)=0$, denoted by $\mathcal{X}^{\#}$ is a Banach space [11] w.r.t. the Lipschitz norm
$\|f\|_{\text{Lip}_{0}}\coloneqq\sup_{x,y\in\mathcal{X},x\neq y}\frac{|f(x)-f(y)|}{\|x-y\|}.$
Let $\mathcal{M}\subseteq\mathcal{X}$ be a subset such that $0\in\mathcal{M}$. Let $A:\mathcal{M}\to\mathcal{X}$ be a Lipschitz map such that $A(0)=0$. Given $x\in\mathcal{M}$ and $f\in\mathcal{X}^{\#}$ satisfying $f(x)=1$, we define two uncertainties of $A$ at $(x,f)\in\mathcal{M}\times\mathcal{X}^{\#}$ as
@@
$\nabla(f,A,x)\coloneqq\|fA-f(Ax)f\|_{\text{Lip}_{0}}.$
@@
$\nabla(f,A,x)\Delta(B,x,f)$ $=\|fA-f(Ax)f\|_{\text{Lip}_{0}}\|Bx-f(Bx)x\|$
@@
$\nabla(f,A,x)\Delta(B,x,f)+\|fB\|_{\text{Lip}_{0}}\Delta(A,x,f)\geq|f([A,B]x)|.$
@@
$\geq|f([A,B]x)|-\|fB\|_{\text{Lip}_{0}}\|Ax-f(Ax)x\|$
$=|f([A,B]x)|-\|fB\|_{\text{Lip}_{0}}\Delta(A,x,f).$
@@
$\nabla(f,A,x)\Delta(B,x,f)+\|fB\|_{\text{Lip}_{0}}\Delta(A,x,-f)\geq|f(\{A,B\}x)|.$
@@
$\geq|f(\{A,B\}x)|-\|fB\|_{\text{Lip}_{0}}\|Ax-(-f)(Ax)x\|$
$=|f(\{A,B\}x)|-\|fB\|_{\text{Lip}_{0}}\Delta(A,x,-f).$
@@
$\nabla(f,A,x)\Delta(B,x,f)+\|fB\|\Delta(A,-x,f)=\nabla(f,A,x)\Delta(B,x,f)+\|fB\|\Delta(A,x,-f)\geq|f(\{A,B\}x)|.$

### v3

@@ 1. Introduction
Let $\mathcal{H}$ be a complex Hilbert space and AA be a possibly unbounded self-adjoint linear operator defined on domain $\mathcal{D}(A)\subseteq\mathcal{H}$. For $h\in\mathcal{D}(A)$ with $\|h\|=1$, define the uncertainty of AA at the point $h$ as
@@
In 1929, Robertson [6] derived the following mathematical form of the uncertainty principle (also known as uncertainty relation) of Heisenberg derived in 1927 [4] (English translation of 1927 original article by Heisenberg). Recall that for two linear operators $A:\mathcal{D}(A)\to\mathcal{H}$ and $B:\mathcal{D}(B)\to\mathcal{H}$, we define $A,B]\coloneqq AB-BA$ and $\{A,B\}\coloneqq AB+BA$.
Theorem $1.$
@@
In this short note, we answer Question 1.3 by deriving an uncertainty principle for Lipschitz maps acting on subsets of Banach spaces. We note that there is a Banach space version of uncertainty principle by Goh and Goodman [3] which differs from the results in this paper. It is interesting to note that the uncertainty principle of Game theory, derived by Székely and Rizzo is nonlinear [8].
@@ 2. Nonlinear Heisenberg-Robertson-Schrodinger Uncertainty Principle
Let $\mathcal{X}$ be a Banach space and $\mathcal{M}\subseteq\mathcal{X}$ be a subset such that $0\in\mathcal{M}$. Recall that the collection of all Lipschitz functions $f:\mathcal{M}\to\mathbb{C}$ satisfying $f(0)=0$, denoted by $\mathcal{M}^{\#}$ is a Banach space [10] w.r.t. the Lipschitz norm
$\|f\|_{\text{Lip}_{0}(\mathcal{M},\mathbb{C})}\coloneqq\sup_{x,y\in\mathcal{M},x\neq y}\frac{|f(x)-f(y)|}{\|x-y\|}.$
Let $\mathcal{M}\subseteq\mathcal{X}$ be a subset such that $0\in\mathcal{M}$. Let $A:\mathcal{M}\to\mathcal{X}$ be a Lipschitz map such that $A(0)=0$. Given $x\in\mathcal{M}$ and $f\in\mathcal{X}^{\#}$ satisfying $f(x)=1$, we define two uncertainties of AA at $(x,f)\in\mathcal{M}\times\mathcal{X}^{\#}$ as
@@
$\nabla(f,A,x)\coloneqq\|fA-f(Ax)f\|_{\text{Lip}_{0}(\mathcal{M},\mathbb{C})}.$
@@
$\nabla(f,A,x)\Delta(B,x,f)$ $=\|fA-f(Ax)f\|_{\text{Lip}_{0}(\mathcal{M},\mathbb{C})}\|Bx-f(Bx)x\|$
@@
$\nabla(f,A,x)\Delta(B,x,f)+\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\Delta(A,x,f)\geq|f([A,B]x)|.$
@@
$\geq|f([A,B]x)|-\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\|Ax-f(Ax)x\|$
$=|f([A,B]x)|-\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\Delta(A,x,f).$
@@
$\nabla(f,A,x)\Delta(B,x,f)+\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\Delta(A,x,-f)\geq|f(\{A,B\}x)|.$
@@
$\geq|f(\{A,B\}x)|-\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\|Ax-(-f)(Ax)x\|$
$=|f(\{A,B\}x)|-\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\Delta(A,x,-f).$
@@
$\nabla(f,A,x)\Delta(B,x,f)+\|fB\|\Delta(A,-x,f)$ $=\nabla(f,A,x)\Delta(B,x,f)+\|fB\|\Delta(A,x,-f)$
$\geq|f(\{A,B\}x)|.$